OpenStudy (anonymous):

a²-b²+6b-9 help my factor this, please

OpenStudy (cwrw238):

group them like this a^2 - 9 + b^2 + 6b now factor first 2 terms and last 2 - this will help

OpenStudy (cwrw238):

thats a^2 - 9 - b^2 + 6b

OpenStudy (anonymous):

(a+3)(a-3)-b(b-6)

OpenStudy (cwrw238):

thats right

OpenStudy (anonymous):

thank you ! i've been trying to figure this out since last night :)

OpenStudy (cwrw238):

you are welcome but that hasn't factored it completely - perhaps that the best we can do - ill check it out

OpenStudy (anonymous):

okay, thanks:)

OpenStudy (cwrw238):

no - hold on this should factor into 2 trinomials =a^2 - b^2 + 6b - 9 = (a + b)(a - b) + 6b - 9 so last term in each parentheses will b + 3 and -3 to give -9 i think its ( a + b - 3)(a - b + 3)

OpenStudy (cwrw238):

lets expand that to check = a^2 -ab + 3a + ab - b^2 + 3b -3a + 3b - 9 = a^2 - b^2 + 6b - 9 yes

OpenStudy (cwrw238):

these are tricky ones

OpenStudy (anonymous):

cant it be (a+b)(a-b)+[3(2b-3) 3(a+b)(a-b)(2b-3)

OpenStudy (cwrw238):

no - that final part does not equal previous line :- (a+b)(a-b) is multiplied by in last but not in previous

OpenStudy (cwrw238):

* multiplied by 3

OpenStudy (cwrw238):

before you can take out part of each term there must be common terms eg 3(x + 2) - y(x+ 2) = (x+2)(3-y)

OpenStudy (cwrw238):

there are no common terms in (a+b)(a-b)+[3(2b-3)

OpenStudy (anonymous):

i think we're allowed to have the final answer in 3 brackets

OpenStudy (anonymous):

i'm very confused, i still dont get why we cant factor out 3 in 6b-9

OpenStudy (cwrw238):

yes you can that = 3(2b - 3) and (a+b)(a - b) + 3(2b - 3) is another way of writing the original function but its not factoring it completely

OpenStudy (cwrw238):

and there are no common factors in the above 2 terms so thats as far as you cab take it

OpenStudy (anonymous):

yeah okay i understand now, thanks :)

OpenStudy (cwrw238):

yw

Directrix (directrix):

@burhan101 This expression factors as shown below: a²-b²+6b-9 = a^2 - ( b^2 -6b + 9) = a^2 - (b - 3) ^2 = [a + (b - 3) ] [a - (b - 3)] = (a + b - 3) (a - b + 3)